I need help I'm studying a test and my teacher gave me thee 8 problems as a review. I know that is is a little much to do 8 problems but I do not no what to do at all then I will just fail the test
Thank u
Btw these are 2x2
Y=2x-10
2y=x-8
Y=-x+4
Y=3x
2y=x+1
-2x-y=7
X+2y=14
Y=3x-14
2x-3y=14
Y=1/2x-7/2
-3x-2y=5.5
X+3y=7.5
6x-4y=54
-9x+2y=-69
Y=-x/2-4
-2x-y=-5
Please answer these questions ASAP I need them for tonight so I can study also answer them in substition

Answer by Alan3354(69443)

Y=2x-10
2y=x-8
Y=-x+4
-------------
y is solved for, so sub for y in the 1st eqn
2y = x - 8
2(x-4) = x-8
2x-8 = x-8
x-8 = -8
x = 0
---
y = 4
===============================
Y=3x
2y=x+1
Same as above
===============================
-2x-y=7
X+2y=14
Solve for x or y in 1 eqn, then sub into the other.
=================
Y=3x-14
2x-3y=14
Same as #1
=================
Y=1/2x-7/2
-3x-2y=5.5
------
X+3y=7.5
6x-4y=54
----------
-9x+2y=-69
Y=-x/2-4
-2x-y=-5